# Momentum-Aware Trajectory Optimisation using Full-Centroidal Dynamics and Implicit Inverse Kinematics

### v2

@@ I-B Related Work
Simplifications, such as linearised centroidal dynamics models, tend to be used to enforce convexity [6]. However, these fail to encapsulate the system’s true nonlinearities, critical for dynamic stability during complex manoeuvres. The authors in [7] represent the robot’s dynamics in their centroidal-form and by using its full kinematics, they directly optimise for the joint trajectories of the actuated degrees of freedom. In comparison, our approach only optimises task-space quantities, offering a more intuitive planning scheme with lower state-space dimensionality, resulting in an efficient formulation and resolution via DDP. Reducing the dimensions of the problem is crucial, since in [2] the classical full-centroidal dynamics model is proven to have the same computational complexity with the Full-Rigid Body Dynamics (FRBD) models in the quadrupeds’ case. Approaches using FRBD models, such as [8] & [2] offer promise but suffer from their high-degree of nonlinearity along with the inherent complexities of planning in the less intuitive joint-space. It is true for [8] that part of the computation can be parallelised, however this requires a specialised DDP solver. Some other innovative approaches have used an alternating optimisation paradigm, attempting dynamic consensus between a reduced-order model and FRBD [9, 10]. However, these approaches do not come without their set of complications. The alternating optimisation cycles, along with the inertia-dependent dynamic consensus constraints impose a performance overhead, threatening the feasibility of real-time application. Moreover, as the time-horizon increases, the growing dimensions of the Karush-Kuhn-Tucker (KKT) matrix challenge general-purpose solvers, leading to higher computational times and problematic convergence.
@@ II-A Problem Formulation
The state vector in Eq. (2) lies on the differentiable manifold of the admissible states () and comprises the Cartesian position and orientation of the base-origin , the corresponding base velocities and each foothold’s 3D-position for the Left-Fore (), Left-Hind (), Right-Fore () and Right-Hind () legs. Moreover, the control-input vector in Eq (2) lies on the Euclidean manifold of admissible controls () which is defined by the box-constraint of Eq. (1), and consists of each foothold’s contact force () and Cartesian velocity (). Note that the state is a Lie-Group [13] and as such, it satisfies different axioms and operations from the Euclidean ones, such as the Integration and Difference () operations [14]. Lastly, it does not include footholds’ velocities, effectively reducing the problem’s dimensions.
@@ II-B Dynamics Model
As previously mentioned, the Full-Centroidal Dynamics model used in this work (see Eq. (II-B)) encapsulate all of the system’s nonlinearities and the nonholonomic constraint of angular momentum. The chosen inertia has no simplifications, while for its calculation, the classic Articulated Body Algorithm (ABA) [15] is not used. The model has a translational part shown in Eq. (3a) and a rotational one in Eq. (3b).
The translational part is simply Newton’s Second-Law, calculating the translational acceleration of the robot’s CoM () in which represents the gravitational acceleration expressed in the World frame (W). On the other hand, the rotational part in Eq. (3b) is not that simple. The calculation of the angular acceleration () not only imposes the nonholonomy to the dynamics, but also major nonlinearities manifest, such as the coupling of the contact forces with the CoM and footholds’ positions and the nonlinear configuration-dependent inertia matrix.
In order to calculate the inertia and CoM, the configuration of the robot’s legs () is necessary. However, including it in the state of the system would have not only been redundant, but also it would have downgraded the performance of the OCP significantly. Augmenting the state with explicit joint-angles, as opposed to their implicit calculation, renders the problem more sparse. Nonetheless, this approach does not mitigate the inherent nonlinearities and the associated local minima of the underlying OC problem, as the dynamics continue to be represented by the same model in both scenarios. Furthermore, such augmentation necessitates an additional four tensor contractions for the computation of the dynamics’ derivatives, attributed to the requirement of differentiating the inverse kinematic Jacobians that map the foothold velocities to the joint ones, as Eq. (4) illustrates.
@@
Ultimately, the dynamics constraint of the OCP in Eq. (1) is composed as in Eq. (6) and lives in the Euclidean tangent-space () of the admissible states. More specifically, the control-input directly affects the footholds, which in turn, along with the contact forces, affect the CoM’s dynamics (see Eq. (II-B)).
@@ II-C Calculation of Derivatives
Box-FDDP is a gradient-based method. Hence, the state and control-input derivatives of the dynamics and cost, along with the cost’s hessians are required. The solver approximates the cost with a linear quadratic function and finds the descend direction at each iteration. The inherent limitation of DDP-based solvers is that they cannot handle state-dependent constraints explicitly, so kinematic reachability and friction-cone constraints must be encoded as penalties, adding complexity and sensitivity to the cost function. In order to have robust convergence and increase the overall performance of the algorithm, an efficient derivation of the fully analytical derivatives and cost-hessians has been performed. This proved to be a challenging task.
Furthermore, our hybrid state-vector does not include the velocities of the footholds, hence a new custom Crocoddyl action model had to be implemented, while for its integration, the library’s symplectic Euler integration has been employed, a simple variational integrator [16] that conserves the system’s geometric properties and phase-space symmetries, ensuring enhanced energy conservation.
@@
Note that the derivative of the Difference operator is non-trivial only for the elements of the state (i.e. ), hence Eq. (9) shows the derivation only for these elements. Note that represent the tangent-space lie-algebra vectors of the manifold [14]. These vectors are related to the base’s position and orientation with the Exponential and Logarithmic mappings. Furthermore, the left-Jacobian () of the quaternion (), along with matrix are derived in [14].
@@
We need a workspace violation avoidance mechanism to avoid kinematic singularities and prevent irregular configurations and derivatives. Each foothold that violates the robot’s kinematic constraint is normalised to be inside the effective workspace of the leg. Then, the modified state is fed into the IK algorithm.
@@ III-C Trajectory Optimisation with Minimal Reference Inputs
Dynamics | Convergence | |||
Model | calc [s] | calcDiff [s] | Jump | Rot. Jump |
@@ III-D Computational Performance
Our condensed formulation results in a a lower-dimensionality than the FRBD equivalent models and the improvements in convergence speed are analysed here. According to Table II, our framework achieves a considerable speedup not only in the integration of the dynamics, but also in the demonstrated convergence compared to methods in literature [3]. Finally, our code-base is not optimised for real-time performance, since no re-planning is used in the current work, hence its true potential is expected to be higher. Both tasks in Table II have 6.03s duration without warm-starting, while the runtime statistics are calculated for 1000 samples. Note that planning such long-horizon trajectories with penalty-based solvers, such as Box-FDDP, is challenging since soft-constraints result in local-minima, compromising convergence.
@@ V Conclusion
The quest for enhanced agility in legged robots, particularly within the realm of trajectory optimisation, remains a challenging endeavour. This work presented a comprehensive trajectory optimisation framework tailored for the quadruped ANYmal C. Specifically, we presented a superior Full-Centroidal Dynamics formulation, compared to the ones found in literature [6, 7, 9], capable of automatic foothold discovery and precise inertia tracking, without using redundant states. To the best of our knowledge, this is the first time that the analytic IK are implicitly incorporated to increase performance without compromising the fidelity of the model. These innovations not only enhance convergence but also provide a more intuitive planning process by optimising task-space quantities with the additional advantage of requiring minimal regularisation. Our approach has been validated through real-world experiments providing definitive answers to our guiding questions as stated in Section III.

### v3

@@ I-B Related Work
Simplifications, such as linearised centroidal dynamics models with constant inertia, tend to be used to enforce convexity [6, 7]. However, these fail to encapsulate the system’s true nonlinearities, critical for dynamic stability during complex manoeuvres. The authors in [8] represent the robot’s dynamics in their centroidal-form and by using its full kinematics, they directly optimise for the joint trajectories of the actuated degrees of freedom. In comparison, our approach only optimises task-space quantities, offering a more intuitive planning scheme with lower state-space dimensionality, resulting in an efficient formulation and resolution via DDP. Reducing the dimensions of the problem is crucial, since in [2] the classical full-centroidal dynamics model is proven to have the same computational complexity with the Full-Rigid Body Dynamics (FRBD) models in the quadrupeds’ case. Approaches using FRBD models, such as [9] & [2] offer promise but suffer from their high-degree of nonlinearity along with the inherent complexities of planning in the less intuitive joint-space. It is true for [9] that part of the computation can be parallelised, however this requires a specialised DDP solver. Some other innovative approaches have used an alternating optimisation paradigm, attempting dynamic consensus between a reduced-order model and FRBD [10, 11]. However, these approaches do not come without their set of complications. The alternating optimisation cycles, along with the inertia-dependent dynamic consensus constraints impose a performance overhead, threatening the feasibility of real-time application. Moreover, as the time-horizon increases, the growing dimensions of the Karush-Kuhn-Tucker (KKT) matrix challenge general-purpose solvers, leading to higher computational times and problematic convergence.
@@ II-A Problem Formulation
The state vector in Eq. (2) lies on the differentiable manifold of the admissible states () and comprises the Cartesian position and orientation of the base-origin , the corresponding base velocities and each foothold’s 3D-position for the Left-Fore (), Left-Hind (), Right-Fore () and Right-Hind () legs. The control-input vector from Eq.(2) resides on the Euclidean manifold of admissible controls (), governed by the box constraint in Eq.(1). It includes each foothold’s contact force () and Cartesian velocity (). During the swing-phase, contact forces are set to zero with free foothold velocities, whereas in the contact-phase, velocities are zero and forces are free, using the box-constraint in Eq. (1). The system state, modelled as a Lie Group [14], follows non-Euclidean operations like Integration and Difference [15], excluding foothold velocities to reduce dimensionality.
@@ II-B Dynamics Model
As previously mentioned, the Full-Centroidal Dynamics model used in this work (see Eq. (3)) encapsulate all of the system’s nonlinearities and the nonholonomic constraint of angular momentum. The chosen inertia has no simplifications, while for its calculation, the classic Articulated Body Algorithm (ABA) [16] is not used. The model has a translational part shown in Eq. (3a) and a rotational one in Eq. (3b).
The translational part is simply Newton’s Second-Law, calculating the translational acceleration of the robot’s CoM () in which represents the gravitational acceleration expressed in the World frame (W). On the other hand, the rotational part in Eq. (3b) is not that simple. The calculation of the angular acceleration () not only imposes the nonholonomy to the dynamics, but also major nonlinearities manifest, such as the coupling of the contact forces with the CoM and footholds’ positions and the nonlinear configuration-dependent inertia matrix () .
In order to calculate the inertia and CoM, the configuration of the robot’s legs () is necessary. However, including in the state of the system would have not only been redundant, but also it would have downgraded the performance of the OCP significantly. Even though, augmenting the state with explicit joint-angles, as opposed to their implicit calculation, renders the problem more sparse, this does not mitigate the inherent nonlinearities and the associated local minima of the underlying OC problem, as the dynamics continue to be represented by the same model in both scenarios. Furthermore, such augmentation necessitates an additional four tensor contractions for the computation of the dynamics’ derivatives, attributed to the requirement of differentiating the inverse kinematic Jacobians that map the foothold velocities to the joint ones, as Eq. (4) illustrates.
@@
Ultimately, the dynamics constraint of the OCP in Eq. (1) is composed as in Eq. (6) and lives in the Euclidean tangent-space () of the admissible states. More specifically, the control-input directly affects the footholds, which in turn, along with the contact forces, affect the CoM’s dynamics (see Eq. (3)).
@@ II-C Calculation of Derivatives
Box-FDDP is a gradient-based method. Hence, the state and control-input derivatives of the dynamics and cost, along with the cost’s hessians are required. The solver approximates the cost with a linear quadratic function and finds the descend direction at each iteration. The inherent limitation of vanilla DDP solvers is that they cannot handle state-dependent constraints explicitly, so kinematic reachability and friction-cone constraints must be encoded as penalties, adding complexity and sensitivity to the cost function. In order to have robust convergence and increase the overall performance of the algorithm, an efficient derivation of the fully analytical derivatives and cost-hessians has been performed. This proved to be a challenging task.
Our hybrid state-vector excludes foothold velocities, requiring the implementation of a custom Crocoddyl action model. For integration, the symplectic Euler method from the library was used, a variational integrator [17] that preserves the system’s geometric properties and phase-space symmetries, ensuring better energy conservation.
@@
Note that the derivative of the Difference operator is non-trivial only for the elements of the state (i.e. ), hence Eq. (LABEL:eq:diff_x) shows the derivation only for these elements. Note that represent the tangent-space lie-algebra vectors of the manifold [15]. These vectors are related to the base’s position and orientation with the Exponential and Logarithmic mappings. Furthermore, the left-Jacobian () of the quaternion (), along with matrix are derived in [15].
@@
A workspace violation avoidance mechanism is needed to prevent kinematic singularities and irregular configurations. Footholds violating kinematic constraints are normalised within the leg’s effective workspace before being passed to the IK algorithm.
@@ III-C Trajectory Optimisation with Minimal Reference Inputs
Dynamics [s] | Convergence [] | |||
Model | calc | calcDiff | Jump | Rot. Jump |
@@ III-D Computational Performance
Our condensed formulation results in a a lower-dimensionality than the FRBD equivalent models and the improvements in convergence speed are analysed here. According to Table II, our framework achieves a considerable speedup not only in the integration of the dynamics, but also in the demonstrated convergence compared to methods in literature [3]. Finally, our code-base is not optimised for real-time performance (hence the higher runtime of our method’s calcDiff() compared to the FRBD’s one), since no re-planning is used in the current work, thus its true potential is expected to be higher. Both tasks in Table II have 6.03s duration without warm-starting, while the runtime statistics are calculated for 1000 samples. Note that planning such long-horizon trajectories with penalty-based solvers, such as Box-FDDP, is challenging since soft-constraints result in local-minima, compromising convergence.
@@ V Conclusion
The quest for enhanced agility in legged robots, particularly within the realm of trajectory optimisation, remains a challenging endeavour. This work presented a comprehensive trajectory optimisation framework tailored for the quadruped ANYmal C. We introduced a superior Full-Centroidal Dynamics formulation, outperforming existing methods[6, 8, 10], with automatic foothold discovery and precise inertia tracking, without redundant states. To the best of our knowledge, this is the first to implicitly incorporate analytic IK for improved performance without compromising model fidelity. Our innovations enhance convergence and simplify planning by optimising task-space quantities with minimal regularisation, validated through real-world experiments, as detailed in Section III.